karissafrazier:

Solve the system of equations and choose the correct ordered pair 4x+7y= 47 5x-4y= -5 A. (3,5) B. (4,5) C. (3,7) D. (4,7)

Astro:

Again, Do you know what to do first?

karissafrazier:

tbh not really- not even my fwb knows how to do this

VIBE:

@VIBE wrote:
Rearrange the equation by subtracting what is to the right of the equal sign from both sides of the equation : 4*x+7*y-(47)=0 Solve 4x+7y-47 = 0 Tiger recognizes that we have here an equation of a straight line. Such an equation is usually written y=mx+b ("y=mx+c" in the UK). "y=mx+b" is the formula of a straight line drawn on Cartesian coordinate system in which "y" is the vertical axis and "x" the horizontal axis. In this formula : y tells us how far up the line goes x tells us how far along m is the Slope or Gradient i.e. how steep the line is b is the Y-intercept i.e. where the line crosses the Y axis The X and Y intercepts and the Slope are called the line properties. We shall now graph the line 4x+7y-47 = 0 and calculate its properties Notice that when x = 0 the value of y is 47/7 so this line "cuts" the y axis at y= 6.71429 y-intercept = 47/7 = 6.71429 When y = 0 the value of x is 47/4 Our line therefore "cuts" the x axis at x=11.75000 x-intercept = 47/4 = 11.75000 Slope is defined as the change in y divided by the change in x. We note that for x=0, the value of y is 6.714 and for x=2.000, the value of y is 5.571. So, for a change of 2.000 in x (The change in x is sometimes referred to as "RUN") we get a change of 5.571 - 6.714 = -1.143 in y. (The change in y is sometimes referred to as "RISE" and the Slope is m = RISE / RUN) Slope = -1.143/2.000 = -0.571 1. Slope = -1.143/2.000 = -0.571 2. x-intercept = 47/4 = 11.75000 3. y-intercept = 47/7 = 6.71429

karissafrazier:

i hate math...............

VIBE:

Why?

karissafrazier:

always so complicated

bigjojo88:

1 Solve for xx in 4x+7y=474x+7y=47. x=\frac{47-7y}{4} x= 4 47−7y ​ 2 Substitute x=\frac{47-7y}{4}x= 4 47−7y ​ into 5x-4y=-55x−4y=−5. \frac{5(47-7y)}{4}-4y=-5 4 5(47−7y) ​ −4y=−5 3 Solve for yy in \frac{5(47-7y)}{4}-4y=-5 4 5(47−7y) ​ −4y=−5. y=5 y=5 4 Substitute y=5y=5 into x=\frac{47-7y}{4}x= 4 47−7y ​ . x=3 x=3 5 Therefore, \begin{aligned}&x=3\\&y=5\end{aligned} ​ x=3 y=5 ​

karissafrazier:

can i get a reply lock on this?